# Identifiability of nonlinear ODE Models with Time-Varying Parameters: the General Analytical Solution and Applications in Viral Dynamics

### v4

@@ Abstract
The paper is organized as follows. Chapter 2 provides a basic mathematical characterization of the problem together with basic mathematical concepts and definitions. It also introduces an elementary and illustrative example from robotics (from now on, the case study). We refer to the case study throughout all the paper to better illustrate all the theoretical concepts and the analytical procedures here introduced. In particular, each section ends with a subsection where the theoretical concepts provided by the section are applied to the case study. Chapter 3 provides a summary of our recent solution of the nonlinear UIO introduced in [2]. Chapter 4 introduces the systematic procedure to obtain the identifiability of any ODE model (i.e., Algorithm 4). Chapter 5 introduces the set of differential equations to generate indistinguishable unknown inputs and initial states, simultaneously (i.e., Equations (5.5) and (5.7)). Chapters 6, 7, and 8 investigate the HIV model discussed in [3, 35, 40, 41, 42], the Covid-19 model introduced in [43, 44], and the genetic toggle switch model discussed in [35]. Chapter 9 provides the study of further very simple examples. Finally, Chapter 10 provides our conclusion.
@@ 2.3 Basic algebraic operations and notions
    $L=m$, $\tau^{0}=g^{0},~{}\tau^{1}=g^{1},\ldots,\tau^{m}=g^{m}$, which are the first $m+1$ vector fields that appear in (2.2) (once the unknown inputs are re-ordered, in accordance with the $\mathcal{R}$ operation, defined in Section 3.1.3), and $\sigma=~{}^{m}\nu$, which is the inverse of ${}^{m}\mu$ defined in Section 3.3.1 (see Equation (3.6)). For simplicity, we denote this operation by []${}^{m}[\cdot]^{\gamma}$ (instead of []${}^{{}^{m}\nu}_{~{}~{}g}[\cdot]^{\gamma}$).
@@ 2.6 Definition of the case study
The solution consists of a recursive algorithm (Algorithm $1$Algorithm 1 uses the automatic procedures provided by Algorithm 2 and Algorithm 3. Often, when we refer to Algorithm 1, we actually mean all the three algorithms (1, 2, and 3).) that, given a system characterized by (2.2), automatically builds the entire observation space in a finite number of iterations. In particular, it builds the so-called observability codistribution, which is, by definition, the span of the gradients of all the observable functions. Algorithm 1 builds the generators of the observability codistribution. To this regard, note that, in the algorithm, statements like ”${\Omega}=\textnormal{span}\left\{\nabla h_{1},\ldots,\nabla h_{p}\right\}$” (e.g., Line 2 of Algorithm 1) require no action. The algorithm builds the codistribution by building a set of its generators. When it recursively updates the codistribution it simply computes its new generators (and this is obtained by applying certain operations directly on the generators of the previous codistribution or by computing new covectors starting from the quantities that define the original system $\Sigma$ in (2.2)). In the following subsections, we provide all the ingredients necessary to implement Algorithm 1.
1:  Set $\Sigma$ the system in (2.2).
2:  $\Omega=\textnormal{span}\left\{\nabla h_{1},\ldots,\nabla h_{p}\right\}$.
3:  boolean Continue=true.
4:  while 1 { This is the MAIN loopdo
5:     if $\textnormal{deg}_{w}\left({\Omega}\right)==m_{w}$ then
6:        break main loop.
7:     end if
8:     $m=\textnormal{deg}_{w}\left({\Omega}\right)-1$.
9:     while $\textnormal{deg}_{w}\left({\Omega}\right)>m$ $\textnormal{deg}_{w}\left({\Omega}\right)\neq m_{w}$ do
10:        $m=\textnormal{deg}_{w}\left({\Omega}\right)$.
11:        $[\widetilde{h}_{1},\ldots,\widetilde{h}_{m}]=\mathcal{S}(\Sigma,~{}\Omega)$,   $\Sigma=\mathcal{R}(\Sigma,~{}\widetilde{h}_{1},\ldots,\widetilde{h}_{m})$.
12:        Compute ${}^{m}\mu$, ${}^{m}\nu$, ${}^{m}\widehat{g}^{\alpha}$ (Eqs (3.6-3.7)).
13:        $\Omega=\left<\left.f^{1},\ldots,f^{m_{u}},\left\{\begin{array}[]{lll}{}^{m}\widehat{g}^{0},&\ldots,&{}^{m}\widehat{g}^{m}\\ g^{m+1},&\ldots,&g^{m_{w}}\\ \end{array}\right\}~{}\right|\Omega\right>$
14:     end while
15:     if $\textnormal{deg}_{w}\left({\Omega}\right)==m_{w}$ then
16:        break main loop.
17:     end if
18:     if $m_{u}\neq 0$ then
19:        Run Algorithms 2 and 3 to compute $\widehat{k}^{m}=s^{m}_{x}+r^{m}$.
20:     else
21:        $\widehat{k}^{m}=0$.
22:     end if
23:     Set $k=0$ and ${\Omega_{0}=\Omega}$
24:     while 1 {This is the NESTED loopdo
25:        $k=k+1$.
26:        Reset $[\Sigma,~{}\Omega_{k-1}]=\mathcal{A}(\Sigma,~{}m,~{}\Omega_{k-1})$.
27:        Compute ${}^{m}\mu$, ${}^{m}\nu$, ${}^{m}\widehat{g}^{\alpha}$ (Eqs (3.6-3.7)).
28:        $\Omega^{*}=\Omega_{k-1}$.
29:        if ($k\leq~{}\widehat{k}^{m}$)&($k\geq 2$then
@@
31:        end if
32:        $\Omega_{k}=\left<\left.f^{1},\ldots,f^{m_{u}},\left\{\begin{array}[]{lll}{}^{m}\widehat{g}^{0},&\ldots,&{}^{m}\widehat{g}^{m}\\ g^{m+1},&\ldots,&g^{m_{w}}\\ \end{array}\right\}~{}\right|\Omega^{*}\right>$.
33:        if $(k-1\geq~{}\widehat{k}^{m})\&(\Omega_{k}==\Omega_{k-1})${ The highest unknown input degree of reconstructability is $m<m_{w}$ (system not canonic), and the observability codistribution is $\Omega_{k}$then
34:           Set $\mathcal{O}=\Omega_{k}$.
35:           Continue=false.
36:           break main and nested loop
37:        end if
38:        if $\textnormal{deg}_{w}\left({\Omega_{k}}\right)>m$ then
39:           Set $\Omega=\Omega_{k}$.
40:           break nested loop
41:        end if
42:     end while
43:  end while
44:  if Continue then
45:     $[\widetilde{h}_{1},\ldots,\widetilde{h}_{m_{w}}]=\mathcal{S}(\Sigma,~{}\Omega)$.
46:     Compute $\mu$, $\nu$, $\widehat{g}^{\alpha}$, and $\widetilde{\mathcal{O}}$ (Eqs (3.3-3.5)).
47:     $\mathcal{O}=\left<f^{1},\ldots,f^{m_{u}},\widehat{g}^{0},\widehat{g}^{1},\ldots,\widehat{g}^{m_{w}}~{}\left|~{}\Omega+\widetilde{\mathcal{O}}\right.\right>$
48:  end if
Algorithm 1 The algorithm that builds the observable codistribution ($\mathcal{O}$) for any nonlinear system characterized by (2.2). Line 19 runs Algorithm 2 and 3. All the functions/operations that appear in the algorithm are defined in Sections 3.1-3.3. Note that, for TV systems, the Lie derivative operator along $\widehat{g}^{0}$ and ${}^{m}\widehat{g}^{0}$, i.e., $\mathcal{L}_{\widehat{g}^{0}}$ and $\mathcal{L}_{{}^{m}\widehat{g}^{0}}$, must be replaced by ˙$\dot{\mathcal{L}}_{\widehat{g}^{0}}:=\mathcal{L}_{\widehat{g}^{0}}+\frac{\partial}{\partial t}$ and ˙$\dot{\mathcal{L}}_{{}^{m}\widehat{g}^{0}}:=\mathcal{L}_{{}^{m}\widehat{g}^{0}}+\frac{\partial}{\partial t}$, respectively.
1:  Set $\Sigma^{\prime}=\Sigma$, $m$, $\widetilde{h}_{1},\ldots,\widetilde{h}_{m}$, as at Line 11 of Algorithm 1.
@@
10:     end if
11:  end loop
Algorithm 2 The algorithm that returns the integer $s^{m}_{x}$ at Line 19 of Algorithm 1.
1:  Set $\Sigma^{\prime}=\Sigma$ and $m$ as at Line 11 of Algorithm 1.
2:  Set $k=0$, ${}^{m}\Delta_{0}=\textnormal{span}\left\{f^{1},\ldots,f^{m_{u}}\right\}$.
@@
5:     Reset $[\Sigma^{\prime},~{}^{m}\Delta_{k-1}]=\mathcal{A}(\Sigma^{\prime},~{}m,~{}^{m}\Delta_{k-1})$
6:     ${}^{m}\Delta_{k}=~{}^{m}\Delta_{k-1}+\sum_{\beta=0}^{m}~{}^{m}\left[~{}^{m}\Delta_{k-1}\right]^{\beta}$
7:     if ${}^{m}\Delta_{k}~{}==~{}^{m}\Delta_{k-1}$ then
@@
9:     end if
10:  end loop
Algorithm 3 The algorithm that returns the integer $r^{m}$ at Line 19 of Algorithm 1.
@@ 3.1.1 $\boldsymbol{\deg_{w}(\Omega)}$ operation
The main loop of Algorithm 1 starts by checking if $\textnormal{deg}_{w}(\Omega)==m_{w}$. If this is the case, the main loop is interrupted. Then, Lines 45-47 of Algorithm 1 are executed and Algorithm 1 ends. For the clarity sake, we devote Section 3.2 to describe the behaviour of Algorithm 1 in this case. Before, we provide the definition of further operations adopted by Algorithm 1.
@@ 3.1.4 $\boldsymbol{\mathcal{A}(\Sigma,~{}m)}$ operation
In Algorithms 1, 2, and 3, we also use this operation by including further outputs and further inputs. Specifically, we consider the following two cases:
@@
We can apply this operation multiple consecutive times (and this is the case of Algorithms 1, 2, and 3, as the operation appears in a loop). The resulting systems still satisfy (2.2) and they are all characterized by $m_{u}$ known inputs and $m_{w}$ unknown inputs. In [2], we called these extended systems Finite Unknown Inputs Extensions. In addition, in [2], we introduced the following definition:
@@ 3.2 Algorithm 1 for systems canonic with respect to the unknown inputs
When a system is canonic with respect to its unknown inputs, Algorithm 1 ends with the execution of Lines 45-47 (the boolean variable Continue remains set to ”true”). The system could be directly in canonic form with respect to its unknown inputs or $\Omega$ was obtained in other parts of the algorithm (e.g., during the execution of the nested loop) and the system has been set in canonical form after the determination of one or more observable functions that are not among the outputs. The execution of Line 45 of Algorithm 1 provides the $m_{w}$ scalar functions $\widetilde{h}_{1},\ldots,\widetilde{h}_{m_{w}}$ ($[\widetilde{h}_{1},\ldots,\widetilde{h}_{m_{w}}]=\mathcal{S}(\Sigma,~{}\Omega)$). Then, the algorithm computes the following quantities: $\mu$, $\nu$, $\widehat{g}^{\alpha}$, and the codistribution $\widetilde{\mathcal{O}}$.
@@ 3.3 Algorithm 1 for systems that are not in canonical form
Let us back to the first part of Algorithm 1, at the beginning of the main loop, and let us consider now the case when $\textnormal{deg}_{w}\left({\Omega}\right)<m_{w}$. The system is not in canonical form with respect to its unknown inputs. Algorithm 1 proceeds in this case with the loop at lines 9-14. The body of this loop is executed at least once.  The set of scalar functions $\widetilde{h}_{1},\ldots,\widetilde{h}_{m}$ are obtained by executing the operation $\mathcal{S}(\Sigma,~{}\Omega)$ defined in Section 3.1.2, and the unknown inputs are reordered accordingly ($\mathcal{R}(\Sigma,~{}\widetilde{h}_{1},\ldots,\widetilde{h}_{m})$, Section 3.1.3). Then, the algorithm computes the following quantities: ${}^{m}\mu$, ${}^{m}\nu$, and ${}^{m}\widehat{g}^{\alpha}$.
3.3.1 $\boldsymbol{{}^{m}\mu}$ and $\boldsymbol{{}^{m}\nu}$
The two-index tensor ${}^{m}\mu$ is defined as follows:
$^{m}\mu^{i}_{j}=\mathcal{L}_{g^{i}}\widetilde{h}_{j},~{}~{}~{}i,~{}j=1,\ldots,m$ (3.6)
${}^{m}\mu^{0}_{0}=1,~{}^{m}\mu^{i}_{0}=0,~{}^{m}\mu^{0}_{i}=\frac{\partial\widetilde{h}_{i}}{\partial t}+\mathcal{L}_{g^{0}}\widetilde{h}_{i},~{}i=1,\ldots,m.$

### v5

@@ Abstract
The paper is organized as follows. Chapter 2 provides a basic mathematical characterization of the problem together with basic mathematical concepts and definitions. It also introduces an elementary and illustrative example from robotics (from now on, the case study). We refer to the case study throughout all the paper to better illustrate all the theoretical concepts and the analytical procedures here introduced. In particular, each section ends with a subsection where the theoretical concepts provided by the section are applied to the case study. Chapter 3 provides a summary of our recent solution of the nonlinear UIO introduced in [2]. Chapter 4 introduces the systematic procedure to obtain the identifiability of any ODE model (i.e., Algorithm 5). Chapter 5 introduces the set of differential equations to generate indistinguishable unknown inputs and initial states, simultaneously (i.e., Equations (5.5) and (5.7)). Chapters 6, 7, and 8 investigate the HIV model discussed in [3, 35, 40, 41, 42], the Covid-19 model introduced in [43, 44], and the genetic toggle switch model discussed in [35]. Finally, Chapter 9 provides our conclusion.
@@ 2.3 Basic algebraic operations and notions
    $L=m$, $\tau^{0}=g^{0},~{}\tau^{1}=g^{1},\ldots,\tau^{m}=g^{m}$, which are the first $m+1$ vector fields that appear in (2.2) (once the unknown inputs are re-ordered, in accordance with the $\mathcal{R}$ operation, defined in Section 3.1.3), and $\sigma=\nu$, which is the inverse of $\mu$ defined in Section 3.3.1 (see Equation (3.6)). For simplicity, also in this case, we denote this operation by $[\cdot]^{\gamma}$ (instead of []${}^{\nu}_{g}[\cdot]^{\gamma}$).
@@ 2.6 Definition of the case study
The solution consists of a recursive procedure (Algorithm 1, which uses Algorithm 2, Algorithm 3 and Algorithm 4111Often, when we refer to Algorithm 1, we actually mean all the four algorithms (1-4).). Given a system characterized by (2.2), Algorithm 1 automatically builds the entire observation space in a finite number of iterations. In particular, it builds the so-called observability codistribution, which is, by definition, the span of the gradients of all the observable functions. Algorithm 1 builds the generators of the observability codistribution. To this regard, note that, in the algorithm, statements like ”${\Omega}=\textnormal{span}\left\{\nabla h_{1},\ldots,\nabla h_{p}\right\}$” (e.g., Line 1 of Algorithm 1) require no action. The algorithm builds the codistribution by building a set of its generators. When it recursively updates the codistribution it simply computes its new generators (and this is obtained by applying certain operations directly on the generators of the previous codistribution or by computing new covectors starting from the quantities that define the original system $\Sigma$ in (2.2)). In the following subsections, we provide all the ingredients necessary to implement Algorithm 1.
Algorithm 1 The algorithm to obtain the state observability in the presence of time varying parameters (or unknown inputs). Comments are in blue.
1:Set $\Sigma$ the system in (2.2), $\Omega=\textnormal{span}\left\{\nabla h_{1},\ldots,\nabla h_{p}\right\}$, and the boolean Continue=True.
2$\triangleright$ This is the MAIN loop
3:     if $\textnormal{deg}_{w}\left({\Omega}\right)==m_{w}$  then break main loop.
4:     $m=\textnormal{deg}_{w}\left({\Omega}\right)$.
5:     $[\widetilde{h}_{1},\ldots,\widetilde{h}_{m}]=\mathcal{S}(\Sigma,~{}\Omega)$,   $\Sigma=\mathcal{R}(\Sigma,~{}\widetilde{h}_{1},\ldots,\widetilde{h}_{m})$.
6:     Compute $\mu$, $\nu$, $\widehat{g}^{\alpha}$ (Eqs (3.6-3.7)).
7:     Run Algorithm 2
8:     if Finish then $\mathcal{O}=\Omega_{*}$, $\mathcal{E}=\Sigma$, Continue=false, and break main loop.
9:     if $m_{u}>0$ then
10:         $\Sigma=\Sigma_{*}$, and Compute $\mu$, $\nu$, $\widehat{g}^{\alpha}$ (Eqs (3.6-3.7)).
11:         Compute $\chi_{*}=\psi_{k_{*}-1}^{i_{*}}$ (Eq (3.11)$\triangleright$ Carried out recursively in $\mathcal{W}$. Then, it is expressed in $\mathcal{V}$.
12:         $\Omega=\Omega+\textnormal{span}\left\{\mathcal{L}_{\chi_{*}}\nabla\widetilde{h}_{q_{*}}\right\}$$\triangleright$ When $k_{*}=1$ $\chi_{*}=f^{i_{*}}$
@@
14:     else
[…]
16:     Reset $[\Sigma,~{}\Omega]=\mathcal{A}^{-}(\Sigma,~{}\Omega)$$\triangleright$ This pass again in $\mathcal{W}$.
17:if Continue then
18:     $[\widetilde{h}_{1},\ldots,\widetilde{h}_{m_{w}}]=\mathcal{S}(\Sigma,~{}\Omega)$.
19:     Compute $\mu$, $\nu$, $\widehat{g}^{\alpha}$ (Eqs (3.3-3.4)), and $\widetilde{\mathcal{O}}$ (Eq. (3.5)).
20:     $\mathcal{O}=\left<f^{1},\ldots,f^{m_{u}},\widehat{g}^{0},\widehat{g}^{1},\ldots,\widehat{g}^{m_{w}}~{}\left|~{}\Omega+\widetilde{\mathcal{O}}\right.\right>$ and set $\mathcal{E}=\Sigma$.
Algorithm 2 This algorithm returns the Boolean Finish. When Finish==True, it also returns $\Omega_{*}$. When Finish==False (and $m$ and $m_{u}$ are both non zero) it also returns $\Sigma_{*}$, $k_{*}$, $i_{*}$, and $q_{*}$.
1:if $(m_{u}==0)|(m==0)$ then
2:     $\Omega_{*}=\left<\left.\widehat{g}^{0},\ldots,\widehat{g}^{m}~{}\right|\Omega\right>$.
3:     if ($\widehat{g}^{m+1}\in\Omega_{*}^{\bot})\And(\widehat{g}^{m+2}\in\Omega_{*}^{\bot})\And\ldots\And(\widehat{g}^{m_{w}}\in\Omega_{*}^{\bot}$then
4:         Finish=True.
5:     else
6:         Finish=False.
7:else
8:     Run Algorithms 3 and 4 to compute $\widehat{k}^{m}=s^{m}_{x}+r^{m}$.
9:     for $k=1,\ldots,\widehat{k}^{m}$ do
[…]
11:              $\Omega=\Omega+\sum_{\alpha_{1}=0}^{m}\ldots\sum_{\alpha_{k-1}=0}^{m}\textnormal{span}\left\{\mathcal{L}_{[f^{i}]^{(\alpha_{1},\ldots,\alpha_{k-1})}}\nabla\widetilde{h}_{q}\right\}$
12:              $\Omega_{*}=\left<\left.f^{1},\ldots,f^{m_{u}},\widehat{g}^{0},\ldots,\widehat{g}^{m}~{}\right|\Omega\right>$.
13:              if ($\widehat{g}^{m+1}\in\Omega_{*}^{\bot})\And(\widehat{g}^{m+2}\in\Omega_{*}^{\bot})\And\ldots\And(\widehat{g}^{m_{w}}\in\Omega_{*}^{\bot}$then
14:                  Compute $m$, $\mu$, $\nu$, $\widehat{g}^{\alpha}$ (Eqs (3.6-3.7)).
15:              else
16:                  Finish=False.
17:                  $\Sigma_{*}=\Sigma$, $k_{*}=k$, $i_{*}=i$, $q_{*}=q$.
18:                  return
19:     Finish=True.
Algorithm 3 The algorithm that returns the integer $s^{m}_{x}$ at Line 7 of Algorithm 2.
1:Set $\Sigma^{\prime}=\Sigma$, $m$, $\widetilde{h}_{1},\ldots,\widetilde{h}_{m}$, as at Line 7 of Algorithm 1.
@@
Algorithm 4 The algorithm that returns the integer $r^{m}$ at Line 7 of Algorithm 2.
1:Set $\Sigma^{\prime}=\Sigma$ and $m$ as at Line 7 of Algorithm 1.
2:Set $k=0$, $\Delta_{0}=\textnormal{span}\left\{f^{1},\ldots,f^{m_{u}}\right\}$.
@@
5:     Reset $[\Sigma^{\prime},~{}\Delta_{k-1}]=\mathcal{A}(\Sigma^{\prime},~{}m,~{}\Delta_{k-1})$
6:     $\Delta_{k}=~{}\Delta_{k-1}+\sum_{\beta=0}^{m}~{}\left[\Delta_{k-1}\right]^{\beta}$
7:     if $\Delta_{k}~{}==~{}\Delta_{k-1}$ then
@@ 3.1.1 $\boldsymbol{\deg_{w}(\Omega)}$ operation
The main loop of Algorithm 1 starts by checking if $\textnormal{deg}_{w}(\Omega)==m_{w}$. If this is the case, the main loop is interrupted. Then, Lines 17-20 of Algorithm 1 are executed and Algorithm 1 ends. For the clarity sake, we devote Section 3.2 to describe the behaviour of Algorithm 1 in this case. Before, we provide the definition of further operations adopted by Algorithm 1.
@@ 3.1.4 $\boldsymbol{\mathcal{A}(\Sigma,~{}m)}$ operation
In Algorithms 1, 3, and 4, we also use this operation by including further outputs and further inputs. Specifically, we consider the following two cases:
@@
We can apply this operation multiple consecutive times (and this is the case of Algorithms 1, 3, and 4, as the operation appears in a loop). The resulting systems still satisfy (2.2) and they are all characterized by $m_{u}$ known inputs and $m_{w}$ unknown inputs. In [2], we called these extended systems Finite Unknown Inputs Extensions. In addition, in [2], we introduced the following definition:
@@
3.1.5 $\boldsymbol{\mathcal{A}^{-}(\Sigma,~{}\Omega)}$ operation
This operation is only executed when, in the iteration of the main loop under execution, the Boolean Finish=False, and the remaining steps of this iteration determines a new observable function that increases the unknown input degree of reconstructability. In this case, the state can be augmented by including the unknown inputs with index larger than $m$ (through the operation $\mathcal{A}$). During this iteration of the main loop, the observable function, which increases the unknown input degree of reconstructability, is added to $\Omega$. Let us denote this function by $\theta$. In general, $\theta$ also depends on the quantities $v_{\alpha}$, defined in (3.9). The first operation executed by $\mathcal{A}^{-}$ is to express $\theta$ only in terms of the unknown inputs and the original state. In other words, all the $v_{\alpha}$ that appear in $\theta$ are expressed in terms of the unknown inputs (this is obtained by using (3.9) and its time derivatives when $\theta$ also depends on the time derivatives of $v_{\alpha}$). Then, all the unknown inputs that appear in $\theta$ are added to the original state. The resulting augmented state defines the new system $\Sigma$.
@@ 3.2 Algorithm 1 for systems canonic with respect to the unknown inputs
When a system is canonic with respect to its unknown inputs, Algorithm 1 ends with the execution of Lines 17-20 (the boolean variable Continue remains set to ”true”). The system could be directly in canonic form with respect to its unknown inputs or $\Omega$ was obtained in other parts of the algorithm  and the system has been set in canonical form after the determination of one or more observable functions that are not among the outputs. The execution of Line 18 of Algorithm 1 provides the $m_{w}$ scalar functions $\widetilde{h}_{1},\ldots,\widetilde{h}_{m_{w}}$ ($[\widetilde{h}_{1},\ldots,\widetilde{h}_{m_{w}}]=\mathcal{S}(\Sigma,~{}\Omega)$). Then, the algorithm computes the following quantities: $\mu$, $\nu$, $\widehat{g}^{\alpha}$, and the codistribution $\widetilde{\mathcal{O}}$.
@@ 3.3 Algorithm 1 for systems that are not in canonical form
Let us back to the first part of Algorithm 1, at the beginning of the main loop, and let us consider now the case when $\textnormal{deg}_{w}\left({\Omega}\right)<m_{w}$. The system is not in canonical form with respect to its unknown inputs. The algorithm continues by computing the set of scalar functions $\widetilde{h}_{1},\ldots,\widetilde{h}_{m}$. They are obtained by executing the operation $\mathcal{S}(\Sigma,~{}\Omega)$ defined in Section 3.1.2, and the unknown inputs are reordered accordingly ($\mathcal{R}(\Sigma,~{}\widetilde{h}_{1},\ldots,\widetilde{h}_{m})$, Section 3.1.3). Then, the algorithm computes the following quantities: $\mu$, $\nu$, and $\widehat{g}^{\alpha}$.
3.3.1 $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$
The two-index tensor $\mu$ is defined as follows:
$\mu^{i}_{j}=\mathcal{L}_{g^{i}}\widetilde{h}_{j},~{}~{}~{}i,~{}j=1,\ldots,m$ (3.6)
$\mu^{0}_{0}=1,~{}\mu^{i}_{0}=0,~{}\mu^{0}_{i}=\frac{\partial\widetilde{h}_{i}}{\partial t}+\mathcal{L}_{g^{0}}\widetilde{h}_{i},~{}i=1,\ldots,m.$